# Report for NSF Workshop on AI for Electronic Design Automation

### v1

@@ I Executive Summary
  3. 3.
    AI toolbox for optimization and design: This section focused on challenges raised to AI from chip design. Existing AI techniques rely heavily on well-maintained offline data, a clearly defined objective function, and an offline evaluation. Key recommendations for NSF include the following. Generalization capability and uncertainty estimation become even more critical in chip design. Reinforcement learning and black-box optimization can lead to more effective and efficient search, which is beyond the traditional prediction tasks in general ML. LLM agents could serve as a very powerful tool by automatically putting long pipelines together and getting end-to-end solutions from data collection, tool calling, to verification.
  4. 4.
    AI for test and verification: This session explored the transformative role of artificial intelligence in test and verification within the domain of EDA. Key discussions included methods for learning robust design representations that generalize across multiple hardware design tasks, enhancing the adaptability and efficiency of AI-driven solutions. A focal point was an LLM-aided hardware design framework which is uniquely enabled by G-QED (Generalized Quick Error Detection), a new verification technique for rapidly identifying logic bugs during chip verification. The workshop also addressed the evolving landscape of software development in the era of LLMs, emphasizing their application in fuzz testing and the novel use of LLMs for generating effective fuzzing programs. Another topic covered was security vulnerabilities of AI-based techniques as they are universally being employed in various testing tasks and their unreliabilities can lead to nefarious system breaches. Finally, presenters showcased AI-based optimization strategies to efficiently solve complex SAT (satisfiability) verification problems, highlighting a path toward more scalable and accurate hardware verification processes. Collectively, the session emphasized the synergy between AI and EDA to enable faster, more intelligent design and design verification flows across the design lifecycle.
@@ II Introduction and Background
In the past few years, domain-specific accelerators (DSAs), such as Google’s Tensor Processing Unit (TPU), have been shown to offer significant performance and energy efficiency over general-purpose CPUs. However, DSAs require deep hardware knowledge to achieve high performance and robustness. Unfortunately, there are far fewer hardware designers than software developers. Leveraging AI techniques to further automate chip design becomes the key to meeting the need of rapid change in software development, which is also critical for democratizing hardware design and creating next-generation energy-efficient hardware. This workshop aims to bring researchers and practitioners from both machine learning and EDA communities to discuss how AI can address challenges in different stages of hardware design, promoting open benchmark datasets and open discussions to revolutionize chip design.
@@
  • LLM-based code analysis for PPA prediction and optimization at all stages of EDA pipeline;
  • Multi-modality ML models (e.g., LLM + GNN) for PPA prediction and optimization at all stages of EDA pipeline;
@@
Based on the three key areas of EDA (AI for physical synthesis and design for manufacturing, AI for high-level and logic-level synthesis, and AI for test and verification), we formed round-table discussions to identify key challenges, SOTA approaches, and future directions. Note that, Theme 3 (AI toolbox for optimization and design) speakers and panelists participated in the other three themes’ discussions.
@@ III-B Summary of Existing Efforts and Showstoppers
  1. 1.
    Reinforcement Learning (RL): Used for macro placement tasks, RL has shown success in reducing wirelength and improving placement quality.
  2. 2.
    Surrogate models for design analysis: Faster and approximate models replace traditional analytical analysis for key design metrics such as timing, power, routability, etc, accelerating optimizations.
  3. 3.
    LLM for rule creation: Large language models (LLMs) are being applied to automate DRC code writing and verification.
  4. 4.
    Generative AI for physical design: Leverage generative AI models trained on optimized data to achieve significant speedup over traditional optimization methods.
  5. 5.
    Gradient-based method for physical design: model physical optimization problems as differentiable problems and optimize with gradient-based methods using deep learning frameworks on GPU.
  6. 6.
    Machine Learning (ML) for placement and routing: RL models improve macro-placement, while predictive tools assist in congestion analysis and timing closure.
  7. 7.
    AI in DFM analysis: Techniques for hotspot detection, process window prediction, and yield optimization enhance design reliability.
  8. 8.
    Simulation acceleration: AI-driven surrogate models provide near-real-time predictions, speeding up simulation processes.
  9. 9.
    Generative AI for analog and RF design: Emerging tools like AnalogCoder and PulseRF enable faster design cycles and optimize performance.
  10. 10.
    AI in early feedback loops: AI-powered tools can provide faster feedback during layout stages, real-time DRC (Design Rule Check) violation detection, supporting informed decision-making and improving productivity.
@@ IV-A Problem Description and Challenges
  • Data dhallenges: AI-driven approaches in HLS and LLS heavily rely on high-quality, labeled data, which is scarce, proprietary, sensitive, and costly to produce. The challenge lies in developing data-sharing mechanisms that protect the proprietary aspects of designs while providing sufficient data to train effective AI models. This balance is crucial for advancing AI capabilities in chip design without compromising intellectual property.
@@ IV-B Summary of Existing Efforts and Showstoppers
One serious challenge of any of these predictive models is to track the changes of the toolchain (e.g. the HLS tool) they plan to model. This problem is addressed by Active-CEM [10] which carries out a task transfer learning scheme that leverages a model-based explorer designed to adapt efficiently to changes in toolchains. This approach optimizes sample efficiency by identifying high-quality design configurations under a new toolchain without requiring extensive re-evaluation. It further refines the methodology by incorporating toolchain-invariant modeling.
Another major challenge arises when the trained model is applied to new designs, as the substantial domain shift often results in unsatisfactory performance. One attempt to address this problem is to use a hierarchical MoE [17], which proposes a more domain-generalizable model structure. Different expert networks can learn to deal with different regions in the representation space, and they can utilize similar patterns between the old kernels and new kernels. The MoE can be applied at basic block level, loop/function level, and the graph level, with the assumption that even for a new kernel, it will share common basic blocks and loop structures with some training examples, although the composition of these basic structures can be different. This approach showed some promising results for adapting to new domains. Nevertheless, more effective and robust methods are needed to transfer learning to new designs.
So far these HLS+DL methods focus on pragma insertion, which is very important in defining the microarchitecture of the chip, and rely on the underlying Merlin compiler [9] to carry out limited program transformation, such as loop strip mining and memory access coalescing. However, more substantial transformation of the original program, e.g., with loop ordering and distribution, may achieve considerably better results, as demonstrated in several non-ML-based approache,s including ScaleHLS [36], HIDA [37], Allo [6], StreamHLS [2], and Sisyphus [28]. However, most of these approaches are limited to affine programs so that some analytical models can be derived for performance prediction. It will be interesting to incorporate these methods and expand the design space of the existing HLS+DL methods to include program transformation for non-affine programs.
@@
Having high-quality training data is important for the success of the approaches in both categories. The HLsyn benchmarks [1] provided 80,000 HLS programs with performance and resource utilization data, which was very helpful for several subsequent research (e.g., [29], [10], [17]). The Chrysalis dataset [34] (symbolizing a phase in the life cycle where a ‘bug’ transforms into a ‘butterfly’) includes an extensive collection of over 1,500 HLS designs, annotated with 19 distinct bug types through LLM-based bug injection techniques. A contest was organized at ICCAD 2024 to collect benchmarks for NL2RTL efforts [38]. More efforts are needed in this direction so that we can reach the training data size of ImageNet scale.
@@ IV-D Future Directions and Recommendations to NSF
  2. 2.
    Carefully engineered AI/ML tools: To help chip designs, the following directions are promising: combining multiple ML models such as GNN and LLM for more comprehensive representations of the circuit; capture design hierarchy and pragma transformations and multi-modality attention; task transfer learning with active learning and domain transfer learning with MoEs, etc. In addition, ML may be useful when (1) an IC is implemented many times after small changes (then we have a large dataset for training); (2) faster prototyping of combinatorial solvers; (3) optimization achieved through Reinforcement Learning, even when the optimization objective is not available in a closed form.
  3. 3.
@@
  • Enhanced “Lint” verification and predict problematic code, possible congestion, and timing problems based on learned experience from previous designs.
@@
  2. 2.
    Promote open framework and platform.
  3. 3.
@@ V-A Problem Description and Challenges
Hardware design verification of chips is a major, perhaps the most significant, design challenge today. Pre-silicon verification is used to detect logic design bugs before chips are manufactured. Several industrial reports highlight significant challenges of existing pre-silicon verification practices (e.g., 80% of design effort may be dedicated to verification, first silicon success lowest since 2004 [11]). The rise of hardware accelerators and multi-chiplet systems makes verification even more challenging. Many design variants and their combinations must now be verified thoroughly and quickly (much quicker and with much nimble design teams, unlike processors). Major verification challenges include: (a) tremendous efforts in crafting design-specific assertions, tests, properties, and full functional specifications, and (b) critical bug escapes despite “100% coverage” highlighting the inadequacy of existing approaches [5]. These challenges get further magnified with AI-generated hardware designs – since AI can generate designs quickly, the relative verification effort increases commensurately.
@@
Bugs that escape pre-silicon verification are detected during post-silicon validation (chips are tested in actual system environments to detect and fix bugs). In addition to logic design bugs, post-silicon validation also targets electrical bugs that are caused by subtle interactions between a design and its electrical state (such as cross-talk, power-supply noise, thermal effects). Post-silicon validation is extremely difficult because the internal signals of a chip are no longer available for controllability and observability unlike pre-silicon verification.
@@
Beyond design bugs, manufacturing defects are also a major concern. Software engineers typically assume that commodity compute chip hardware — e.g., CPU, GPU, ML accelerator — always works correctly during its lifetime. Large-scale distributed systems in data centers often assume that hardware errors, if any, immediately cause software execution to stop via crashes and hangs, for example – the fail stop failure assumption. Unfortunately, none of these assumptions holds today. Too many defective compute chips – i.e., with manufacturing defects – are escaping today’s manufacturing tests – at least an order of magnitude more than industrial targets across all compute chip types in data centers. Silent data corruptions (SDCs) caused by test escapes, when left unaddressed, pose a major threat to reliable computing [22].
@@
Reliability failures, largely benign in the past – radiation-induced soft errors, early-life failures, circuit aging – cause unexpected data corruption and expensive system downtimes with severe consequences [7]. Business-as-usual manufacture time testing, large speed/voltage margins, or expensive redundancy for fault-tolerance aren’t scalable or cost-effective anymore.
@@
The advent of AI-based optimization has dramatically reshaped the landscape of hardware security, presenting unprecedented challenges. The attack surface has expanded to an alarming degree, encompassing not only traditional vulnerabilities within the hardware itself but also novel avenues introduced by the AI integration. Each stage of the design and optimization process, now heavily influenced by AI algorithms, necessitates a rigorously defined formal attack model to effectively identify and mitigate potential threats. This is particularly crucial as the very intelligence intended to optimize design and performance simultaneously introduces new vectors for sophisticated attacks, demanding a proactive and comprehensive re-evaluation of security paradigms. Similarly, the robustness of software systems, which use AI at its core, is becoming highly challenging as well. This is featured by the deep system stack from chips to code, as well as the heterogeneous compute involving non-determinism and complex interactions across hardware platforms, runtime environments, and software ecosystems.
@@ V-B Summary of Existing Efforts and Showstoppers
AI techniques for design verification and bug fixing (e.g., LLM- or RL-based such as LLM4SVA discussed in Sec. IV and others) are in their infancy today – especially for “difficult” corner-case bugs that get detected by specific input sequences. Existing verification techniques are already highly challenging and critical bugs escape existing verification flows that are highly time-consuming and resource-intensive in terms of human expertise. Thus, targeting AI techniques that match the quality of existing verification flows isn’t sufficient to ensure design quality and overall design productivity moving forward. Even when bugs are successfully detected, fixing those bugs must also be automated for AI-driven design flows for overall design productivity. Very little progress has been made in the above areas. Needless to say that high-quality training data is crucial for success (e.g., see the discussion on the Chrysalis dataset and corresponding bug injection techniques in Sec. IV).
@@ V-C Opportunities for AI/EDA Collaboration
One of the biggest AI/ML opportunities for chip design is in “creating” highly efficient and robust digital designs with drastic design productivity improvements – e.g., a few days vs. several months. Since verification is the biggest bottleneck in design today, design and verification cannot be treated separately for overall productivity benefits. This creates a unique opportunity to combine AI/ML techniques synergistically with recent breakthroughs in design verification (such as G-QED [5] and its extensions), which includes non-AI/ML techniques, for an overall AI-driven approach (e.g., AI-Boosted Chip Design or ABCD approach in [23]). The key idea is to effectively combine these techniques by exploiting their strengths while avoiding their weaknesses. To achieve this objective, several deeply technical questions must be answered. A few examples include: 1. How can AI/ML techniques be used to create designs with drastically fewer bugs to start with (before any design verification techniques are applied)? Today’s approaches based on naive natural language prompts create designs riddled with bugs despite small design sizes and no advanced architecture features. 2. How can we verify complex AI/ML-generated designs with minimal human intervention using ideas in [5] and others? What verification guarantees can be provided especially for classes of bugs introduced by AI/ML? 3. How can we feed verification results (e.g., counterexamples or bug traces) back to AI/ML for bug fixes? Can we rely on AI/ML-assisted debugging? 4. How do we ensure that bug fixes are correct and they don’t significantly degrade performance and energy efficiency of resulting designs? Can AI/ML and non-AI/ML techniques collaborate on this aspect? 5. How do we ensure that AI/ML understands various architecture features “correctly” and uses them in “appropriate” ways when creating designs?

### v3

@@ I Executive Summary
  • AI toolbox for optimization and design: This section focused on challenges raised to AI from chip design. Existing AI techniques rely heavily on well-maintained offline data, a clearly defined objective function, and an offline evaluation. Key recommendations for NSF include the following. Generalization capability and uncertainty estimation become even more critical in chip design. Reinforcement learning and black-box optimization can lead to more effective and efficient search, which is beyond the traditional prediction tasks in general ML. LLM agents could serve as a very powerful tool by automatically putting long pipelines together and getting end-to-end solutions from data collection and tool calling to verification.
  • AI for test and verification: This session explored the transformative role of artificial intelligence in test and verification within the domain of EDA. Key discussions included methods for learning robust design representations that generalize across multiple hardware design tasks, enhancing the adaptability and efficiency of AI-driven solutions. A focal point was an LLM-aided hardware design framework, which is uniquely enabled by G-QED (Generalized Quick Error Detection), a new verification technique for rapidly identifying logic bugs during chip verification. The workshop also addressed the evolving landscape of software development in the era of LLMs, emphasizing their application in fuzz testing and the novel use of LLMs for generating effective fuzzing programs. Another topic covered was security vulnerabilities of AI-based techniques, as they are universally being employed in various testing tasks, and their unreliability can lead to nefarious system breaches. Finally, presenters showcased AI-based optimization strategies to efficiently solve complex SAT (satisfiability) verification problems, highlighting a path toward more scalable and accurate hardware verification processes. Collectively, the session emphasized the synergy between AI and EDA to enable faster, more intelligent design and design verification flows across the design lifecycle.
@@ II Introduction and Background
In the past few years, domain-specific accelerators (DSAs), such as Google’s Tensor Processing Unit (TPU), have been shown to offer significant performance and energy efficiency over general-purpose CPUs. However, DSAs require deep hardware knowledge to achieve high performance and robustness. Unfortunately, there are far fewer hardware designers than software developers. Leveraging AI techniques to further automate chip design becomes the key to meeting the need for rapid change in software development, which is also critical for democratizing hardware design and creating next-generation energy-efficient hardware. This workshop aims to bring researchers and practitioners from both machine learning and EDA communities to discuss how AI can address challenges in different stages of hardware design, promoting open benchmark datasets and open discussions to revolutionize chip design.
@@
  • LLM-based code analysis for PPA prediction and optimization at all stages of the EDA pipeline;
  • Multi-modality ML models (e.g., LLM + GNN) for PPA prediction and optimization at all stages of the EDA pipeline;
@@
Based on the three key areas of EDA (AI for physical synthesis and design for manufacturing, AI for high-level and logic-level synthesis, and AI for test and verification), we formed round-table discussions to identify key challenges, SOTA approaches, and future directions. Note that Theme 3 (AI toolbox for optimization and design) speakers and panelists participated in the other three themes’ discussions.
@@ III-B Summary of Existing Efforts and Showstoppers
  • Reinforcement Learning (RL): Used for macro placement tasks [33], RL has shown some success in improving placement quality. RL has also been shown for analog transistor sizing to meet tough design specifications [4].
  • Surrogate models for design analysis: Faster and approximate models replace traditional analytical analysis for key design metrics such as timing, power, routability, etc., accelerating optimizations [20].
  • LLM for rule creation: Large language models (LLMs) are being applied to automate DRC code writing and verification [6].
  • Generative AI for physical design: Leverage generative AI models trained on optimized data to achieve a significant boost over traditional optimization methods, e.g., [31].
  • Gradient-based method for physical design: model physical optimization problems as differentiable problems and optimize with gradient-based methods using deep learning frameworks on GPU [27, 30].
  • Machine Learning (ML) for placement and routing, e.g., predictive tools assist in congestion analysis and timing closure [48].
  • AI in DFM analysis: Techniques for hotspot detection [13], printed image [51] and process window predictions, and yield optimization to enhance design robustness.
  • Simulation acceleration: AI-driven surrogate models provide near-real-time predictions [19], speeding up simulation processes.
  • Generative AI for analog and RF design: Emerging tools like AnalogCoder [24] and PulseRF [5] enable faster design cycles and optimize performance.
  • AI in early feedback loops: AI-powered tools can provide faster feedback during layout stages, real-time DRC (Design Rule Check) violation detection [16], supporting informed decision-making and improving productivity.
@@ IV-A Problem Description and Challenges
  • Data challenges: AI-driven approaches in HLS and LLS heavily rely on high-quality, labeled data, which is scarce, proprietary, sensitive, and costly to produce. The challenge lies in developing data-sharing mechanisms that protect the proprietary aspects of designs while providing sufficient data to train effective AI models. This balance is crucial for advancing AI capabilities in chip design without compromising intellectual property.
@@ IV-B Summary of Existing Efforts and Showstoppers
One serious challenge of any of these predictive models is to track the changes of the toolchain (e.g., the HLS tool) they plan to model. This problem is addressed by Active-CEM [14], which carries out a task transfer learning scheme that leverages a model-based explorer designed to adapt efficiently to changes in toolchains. This approach optimizes sample efficiency by identifying high-quality design configurations under a new toolchain without requiring extensive re-evaluation. It further refines the methodology by incorporating toolchain-invariant modeling.
Another major challenge arises when the trained model is applied to new designs, as the substantial domain shift often results in unsatisfactory performance. One attempt to address this problem is to use a hierarchical MoE [25], which proposes a more domain-generalizable model structure. Different expert networks can learn to deal with different regions in the representation space, and they can utilize similar patterns between the old kernels and new kernels. The MoE can be applied at the basic block level, loop/function level, and the graph level, with the assumption that even for a new kernel, it will share common basic blocks and loop structures with some training examples, although the composition of these basic structures can be different. This approach showed some promising results for adapting to new domains. Nevertheless, more effective and robust methods are needed to transfer learning to new designs.
So far, these HLS+DL methods focus on pragma insertion, which is very important in defining the microarchitecture of the chip, and rely on the underlying Merlin compiler [12] to carry out limited program transformation, such as loop strip mining and memory access coalescing. However, more substantial transformation of the original program, e.g., with loop ordering and distribution, may achieve considerably better results, as demonstrated in several non-ML-based approaches, including ScaleHLS [49], HIDA [50], Allo [9], StreamHLS [2], and Sisyphus [40]. However, most of these approaches are limited to affine programs, so that some analytical models can be derived for performance prediction. It will be interesting to incorporate these methods and expand the design space of the existing HLS+DL methods to include program transformation for non-affine programs.
@@
Having high-quality training data is important for the success of the approaches in both categories. The HLsyn benchmarks [1] provided 80,000 HLS programs with performance and resource utilization data, which was very helpful for several subsequent research (e.g., [41, 14, 25]). The Chrysalis dataset [46] (symbolizing a phase in the life cycle where a ‘bug’ transforms into a ‘butterfly’) includes an extensive collection of over 1,500 HLS designs, annotated with 19 distinct bug types through LLM-based bug injection techniques. A contest was organized at ICCAD 2024 to collect benchmarks for NL2RTL efforts [52]. More efforts are needed in this direction so that we can reach the training data size of ImageNet scale.
@@ IV-D Future Directions and Recommendations to NSF
  • Carefully engineered AI/ML tools: To help chip designs, the following directions are promising: combining multiple ML models, such as GNN and LLM, for more comprehensive representations of the circuit; capturing design hierarchy and pragma transformations and multi-modality attention; task transfer learning with active learning and domain transfer learning with MoEs, etc. In addition, ML may be useful when (1) an IC is implemented many times after small changes (then we have a large dataset for training); (2) faster prototyping of combinatorial solvers; (3) optimization achieved through Reinforcement Learning, even when the optimization objective is not available in a closed form.
@@
  • Enhance “Lint” verification and predict problematic code, possible congestion, and timing problems based on learned experience from previous designs.
@@
  • Promote open frameworks and platforms.
@@ V-A Problem Description and Challenges
Hardware design verification of chips is a major, perhaps the most significant, design challenge today. Pre-silicon verification is used to detect logic design bugs before chips are manufactured. Several industrial reports highlight significant challenges of existing pre-silicon verification practices (e.g., 80% of design effort may be dedicated to verification, first silicon success lowest since 2004 [15]). The rise of hardware accelerators and multi-chiplet systems makes verification even more challenging. Many design variants and their combinations must now be verified thoroughly and quickly (much quicker and with much more nimble design teams, unlike processors). Major verification challenges include: (a) tremendous efforts in crafting design-specific assertions, tests, properties, and full functional specifications, and (b) critical bug escapes despite “100% coverage” highlighting the inadequacy of existing approaches [8]. These challenges get further magnified with AI-generated hardware designs – since AI can generate designs quickly, the relative verification effort increases commensurately.
@@
Bugs that escape pre-silicon verification are detected during post-silicon validation (chips are tested in actual system environments to detect and fix bugs). In addition to logic design bugs, post-silicon validation also targets electrical bugs that are caused by subtle interactions between a design and its electrical state (such as cross-talk, power-supply noise, thermal effects). Post-silicon validation is extremely difficult because the internal signals of a chip are no longer available for controllability and observability, unlike pre-silicon verification.
@@
Beyond design bugs, manufacturing defects are also a major concern. Software engineers typically assume that commodity compute chip hardware — e.g., CPU, GPU, ML accelerator — always works correctly during its lifetime. Large-scale distributed systems in data centers often assume that hardware errors, if any, immediately cause software execution to stop via crashes and hangs, for example, the fail-stop failure assumption. Unfortunately, none of these assumptions holds today. Too many defective compute chips – i.e., with manufacturing defects – are escaping today’s manufacturing tests – at least an order of magnitude more than industrial targets across all compute chip types in data centers. Silent data corruptions (SDCs) caused by test escapes, when left unaddressed, pose a major threat to reliable computing [34].
@@
Reliability failures, largely benign in the past – radiation-induced soft errors, early-life failures, circuit aging – cause unexpected data corruption and expensive system downtimes with severe consequences [10]. Business-as-usual manufacturing-time testing, large speed/voltage margins, or expensive redundancy for fault-tolerance aren’t scalable or cost-effective anymore.
@@
The advent of AI-based optimization has dramatically reshaped the landscape of hardware security, presenting unprecedented challenges. The attack surface has expanded to an alarming degree, encompassing not only traditional vulnerabilities within the hardware itself but also novel avenues introduced by the AI integration. Each stage of the design and optimization process, now heavily influenced by AI algorithms, necessitates a rigorously defined formal attack model to effectively identify and mitigate potential threats. This is particularly crucial as the very intelligence intended to optimize design and performance simultaneously introduces new vectors for sophisticated attacks, demanding a proactive and comprehensive re-evaluation of security paradigms. Similarly, the robustness of software systems, which use AI at their core, is becoming highly challenging as well. This is featured by the deep system stack from chips to code, as well as the heterogeneous compute involving non-determinism and complex interactions across hardware platforms, runtime environments, and software ecosystems.
@@ V-B Summary of Existing Efforts and Showstoppers
AI techniques for design verification and bug fixing (e.g., LLM- or RL-based, such as LLM4SVA discussed in Sec. IV and others) are in their infancy today – especially for “difficult” corner-case bugs that get detected by specific input sequences. Existing verification techniques are already highly challenging, and critical bugs escape existing verification flows that are highly time-consuming and resource-intensive in terms of human expertise. Thus, targeting AI techniques that match the quality of existing verification flows isn’t sufficient to ensure design quality and overall design productivity moving forward. Even when bugs are successfully detected, fixing those bugs must also be automated for AI-driven design flows for overall design productivity. Very little progress has been made in the above areas. Needless to say, high-quality training data is crucial for success (e.g., see the discussion on the Chrysalis dataset and corresponding bug injection techniques in Sec. IV).
@@ V-C Opportunities for AI/EDA Collaboration
One of the biggest AI/ML opportunities for chip design is in “creating” highly efficient and robust digital designs with drastic design productivity improvements – e.g., a few days vs. several months. Since verification is the biggest bottleneck in design today, design and verification cannot be treated separately for overall productivity benefits. This creates a unique opportunity to combine AI/ML techniques synergistically with recent breakthroughs in design verification (such as G-QED [8] and its extensions), which includes non-AI/ML techniques, for an overall AI-driven approach (e.g., AI-Boosted Chip Design or ABCD approach in [35]). The key idea is to effectively combine these techniques by exploiting their strengths while avoiding their weaknesses.
To achieve this objective, several deeply technical questions must be answered. A few examples include: 1. How can AI/ML techniques be used to create designs with drastically fewer bugs to start with (before any design verification techniques are applied)? Today’s approaches based on naive natural language prompts create designs riddled with bugs despite small design sizes and no advanced architecture features. 2. How can we verify complex AI/ML-generated designs with minimal human intervention using ideas in [8] and others? What verification guarantees can be provided, especially for classes of bugs introduced by AI/ML? 3. How can we feed verification results (e.g., counterexamples or bug traces) back to AI/ML for bug fixes? Can we rely on AI/ML-assisted debugging? 4. How do we ensure that bug fixes are correct and don’t significantly degrade the performance and energy efficiency of resulting designs? Can AI/ML and non-AI/ML techniques collaborate on this aspect? 5. How do we ensure that AI/ML understands various architecture features “correctly” and uses them in “appropriate” ways when creating designs?